# Joule-Thomson expansion for quantum corrected AdS-Reissner-Nördstrom black holes in Kiselev spacetime with Barrow fractal entropy

### v1

@@
Abstract
How can we detect the difference in the effects of the quantum corrections included in the metric of a spacetime and the quantum corrections included in the entropy of such a system? Recently, J. Barrow designed an expression based directly on black hole (BH) entropy of Bekenstein-Hawking where the geometry of the event horizon can also have an intricate, non smooth, structure, a fractal geometry. These fractal features are represented by a numerical constant parameter, the fractal parameter (FP). Since then, several interesting issues have been explored in the literature. In this work, we investigate the inversion temperature connected to the Joule-Thomson expansion from the thermodynamics of AdS-Reissner-Nördstrom BH by using the Barrow entropy equation where the FP has several values within a certain validity interval. We include quantum corrections in a cosmological fluid that can describe phantom dark matter or quintessence, both in a Kiselev scenario. The description of such physical systems also involves numerical solutions concerning the FP. The results are shown by temperature-pressure curves for multiple values of the parameters of the system used here. In conclusion of our analysis, we also show isenthalpic curves corresponding to fixed-mass BH processes, and we respond numerically to the question made in the first line of this abstract.
@@ I Introduction
The main objective of this paper is to investigate the JT effect modified by the Barrow entropy on quantum corrections and cosmological fluids described by Kiselev spacetime, such as phantom dark energy and quintessence, in an AdS-Reissner-Nördstrom (AdS-RN) BH. An investigation of JTE in a quantum corrected AdS-Reissner-Nordstrom black hole in Kiselev spacetime with standard BeH entropy was presented in [59], so the present work can be considered as an extension to study the effects of the fractal Barrow entropy on JTE isenthalpics and inversion temperature in these BHs.
@@ II Joule-Thomson expansion: a brief review
$\mu_{JT}\,=\,\Big(\frac{\partial T}{\partial P}\Big)_{H}\,$, (2)
which is called the JT coefficient. The sign of $\mu_{JT}$ rules the cooling or heating. As pressure decreases, the variation of pressure is negative, but the variation of temperature may be negative or positive. If the temperature variation is negative (positive), then the JT coefficient is positive (negative). Hence, the gas cools (warms).
@@
$\mu_{JT}\,=\,\Big(\frac{\partial T}{\partial P}\Big)_{H}\,=\,\frac{1}{C_{P}}\,\Bigg[T\Big(\frac{\partial V}{\partial T}\Big)_{P}\,-\,V\Bigg]\,.$ (12)
From Eq. (2) we see that the JT coefficient measures the temperature variation of a gas or liquid when it expands without carrying out work keeping the enthalpy constant. At the inversion temperature, there is no cooling or warming and so there is no temperature variation, and the JT coefficient is zero. At $T_{i}$ the intermolecular forces are in equilibrium. We will have cooling only at $T$ lower than $T_{i}$.
The inversion temperature is characterized by a zero $\mu_{JT}$, so
$T_{i}\,=\,V\,\Big(\frac{\partial T}{\partial V}\Big)_{P}\,\,.$ (13)
@@ III.1 Kiselev black hole
where $d\Omega^{2}\,=\,d\theta^{2}\,+\,\sin^{2}\theta d\phi^{2}$ and
@@ III.2 First law, Barrow entropy, and Inversion Temperature
where the enthalpy $H$ is identified with the BH mass $M=H$. For the inversion temperature, $T_{i}$, it can be shown, as we mentioned in the last section, that $\mu_{JT}=0$.
@@
where $\Delta=0$ recover the Bekenstein-Hawking entropy [3, 4] and $\Delta=1$ means that we have the most intricate geometry, the fractal one, i.e., $0\leq\Delta\leq 1$. Hence, we have a fractal event horizon embedded in the Kiselev spacetime. As we mentioned before, now we have the quantum correction manifested in the entropy expression. In this Barrow model, it also deforms the structure of the event horizon as a function of the $\Delta$-parameter. In addition, it also deforms the way information is stored in the horizon, creating a quantum microstructure or a fractal one. Hence, we expected that the difference between the metric and entropy quantum corrections would show itself in some manner.
@@
Hence, we can write the Hawking temperature as
$T=\frac{1}{2(2+\Delta)(\sqrt{\pi}r_{\!\!{+}})^{\Delta}}\bigg(8Pr_{\!\!{+}}-\frac{Q^{2}}{\pi r_{\!\!{+}}^{3}}+\frac{1}{\pi\sqrt{r_{\!\!{+}}^{2}-a^{2}}}+\frac{3\omega c}{\pi r_{\!\!{+}}^{2+3\omega}}\bigg)\,.$ (21)
Besides, the Joule-Thomson inversion temperature is
$T_{i}=V\bigg(\frac{\partial T}{\partial V}\bigg)_{P}\,\,$, (22)
@@
$T_{i}=\frac{4\pi r_{\!\!{+}}^{3}}{3}\bigg(\frac{\partial T}{\partial r_{\!\!{+}}}\frac{\partial r_{\!\!{+}}}{\partial V}\bigg)_{P}.$ (24)
@@
$T_{i}=\frac{1}{6(2+\Delta)(\sqrt{\pi}r_{\!\!{+}})^{\Delta}}\bigg[8P_{i}(1-\Delta)r_{\!\!{+}}+\frac{Q^{2}(3+\Delta)}{\pi r_{\!\!{+}}^{3}}-\frac{3\omega c(2+\Delta+3\omega)}{\pi r_{\!\!{+}}^{2+3\omega}}+\frac{\Delta a^{2}-(1+\Delta)r_{\!\!{+}}^{2}}{\pi(r_{\!\!{+}}^{2}-a^{2})^{3/2}}\bigg]\,.$ (25)
Using $T=T_{i}$ and $P=P_{i}$ in Eq. (21) and if we subtract this equation from Eq. (25), we obtain a relation between $r_{\!\!{+}}$ and $P_{i}$, which is given by
$\frac{r_{\!\!{+}}^{3}+r_{\!\!{+}}(3+\Delta)(r_{\!\!{+}}^{2}-a^{2})}{(r_{\!\!{+}}^{2}-a^{2})^{3/2}}+8\,P_{i}\,\pi\,r_{\!\!{+}}^{2}(2+\Delta)-\frac{Q^{2}(6+\Delta)}{r_{\!\!{+}}^{2}}+3\,c\,r_{\!\!{+}}^{-1-3\omega}\,\omega\,(5+3\omega+\Delta)=0\,$, (26)
and we have to solve this equation numerically to find $r_{\!\!{+}}(P)$. We can substitute this result into Eq. (25) to construct a $T\times P$ plot where we can see the inversion temperature curve. This will be analyzed in detail in the next section.
@@ IV Numerical results
From now on, we will discuss what happens when we compute the variation of $\Delta$ in the face of the other parameters. First, we will consider the case of fixed parameters $a=0.1,\,c=0.1,\,\omega=-1.0,\,Q=1.0$ and we will vary $\Delta$ with a difference $1/4$ in the interval $\Delta\in[0,1]$. Fig. 1 shows the results of this analysis. In the upper panels, we plot the inversion temperature, Eq. (25), against the pressure $P$ in small and large ranges. In these panels, one sees that increasing $\Delta$ reduces the inversion temperature $T_{i}$, Eq. (25), for fixed $P$, or equivalently, increases $P$ for fixed $T_{i}$. This comes with a change of slope in the curves, without altering the zeros of these functions. Actually, the five curves corresponding to the values of $\Delta$ in the interval $[0,1]$, have the same zero, which is very close to $P_{0}\approx 0.001$. The results of the left panel of Fig. 1 about the decrease of the slope of the inversion temperature curves with increasing values of $\Delta$ are compatible with those found in Fig. 16 of [76] with $F(R)$ gravity, with Figs. 3 and 4 of [77] within Rastall gravity, and with Fig. 4 of [78] within spacetime foam effects.
@@
Then, in the lower panels of Fig. 1, we show the behavior of these functions in the ranges $7\leq P\leq 12$ and $10\leq P\leq 15$, to emphasize the crossing of the inversion temperature curves corresponding to different $\Delta$ values. This means that increasing the fractal dimension $\Delta$ implies decreasing the inversion temperature $T_{i}$ for pressures lower than a certain critical value $P_{c}$, and the opposite behavior above this value. In particular, in the lower left panel of Fig. 1, it is seen that the curves for $\Delta=0$ and $\Delta=1$ cross at $P_{c}\approx 10.2$, which can be determined numerically from Eq. (25) using the values $a=0.1,\,c=0.1,\,\omega=-1.0,\,Q=1.0$, for $\Delta=0$ and $\Delta=1$.
@@
Figure 1: Inversion temperature as a function of the pressure for fixed $\omega=-1.0$, $c=0.1$, $a=0.1$, $Q=1.0$, and the corrections to entropy: $\Delta=0$ (blue line), $\Delta=1/4$ (red line), $\Delta=1/2$ (black line), $\Delta=3/4$ (cyan line) and $\Delta=1$ (green line). The upper panels show small and large behaviors of $T_{i}\times P$, while the lower panels depict different ranges of these functions, with the purpose of showing the intercept points of these curves for different values of $\Delta$.
@@
Figure 2: Inversion temperature as a function of the pressure for fixed $\omega=-1.0$, $c=0.1$, $a=0.1$, small range, and the values for charge: $Q=1.0$ (green line), $Q=2.0$ (red line), $Q=3.0$ (black line) and $Q=6.0$ (blue line). Solid lines are obtained from $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
In Fig. 2, we show the results for $T_{i}\times P$ on both small and large scales, using $\Delta=0$, $\Delta=1/2$, and $\Delta=1$, keeping $a=0.1,\,c=0.1,\,\omega=-1.0$ fixed and varying the charge $Q$ from 1 to 6. Note that increasing $\Delta$ diminishes the slope of the curves $T_{i}\times P$, as already found in Fig. 1 above. In this picture, we also see that the effect on $T_{i}$ of increasing the charge $Q$ is similar to the decrease in the fractal parameter of Barrow $\Delta$, so that from the thermodynamical point of view, the charge $Q$ and $\Delta$ have opposite (inverse) behaviors. Note also that the zeros of these curves are different for each charge value $Q$, corresponding to different initial pressures, and we see that these zeros are not modified by changing $\Delta$. For the small range (left panel), one clearly sees a distinction from the curves with $\Delta=0$, $1/2$, and $1$, while in the large range (right panel) these curves get mixed, since for $\Delta=1$, the slope is almost constant, whilst for $\Delta=0$, and $1/2$ they vary significantly, with decreasing slope with increasing pressure. The results of the right panel of Fig. 2 can be compared with those found in Figs. 3 and 4 of [77] within Rastall gravity.
@@
Figure 3: Inversion temperature as a function of the pressure for fixed $Q=2.5$, $c=0.1$, $a=0.1$, varying $\omega$: $\omega=-1.6$ (green line), $\omega=-2/3$ (red line), $\omega=-1$ (black line) and $\omega=-4/3$ (blue line). Solid lines are obtained from $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
In Fig. 3, we plot $T_{i}\times P$ for fixed $Q=2.5$, $c=0.1$, $a=0.1$, varying the Kiselev parameter $\omega=-1.6$, 2/3-2/3, 1-1, and 4/3-4/3, for $\Delta=0$, 1/2 and 1. For fixed $\Delta$, the slope of the curves for these values of $\omega$ is very similar, although the zeros of the curves are clearly distinguishable. The slopes of these curves are very sensitive to the variation of $\Delta$, while the zeros are not. Again, increasing $\Delta$ decreases the slope of the curves $T_{i}\times P$. In both the small range (left panel) and the large range (right panel), one easily sees a distinction between the curves with different values of $\Delta$, especially in the large range, where the curves with different values of $\omega$ seem to collapse mainly depending on $\Delta$.
@@
Figure 4: Inversion temperature as a function of the pressure for fixed $\omega=-1.0$, $c=0.1$, $Q=1.0$, and the values for the quantum corrections: $a=0.1$ (blue line), $a=0.3$ (red line), $a=0.4$ (black line). Solid lines are obtained with $\Delta=0$, dotted with $\Delta=1/2$, and dashed lines with $\Delta=1$.
@@
To study the interplay between the effects of variation of $\Delta$ and the quantum correction parameter $a$, we present in Fig. 4 the graph $T_{i}\times P$ for fixed $\omega=-1.0$, $c=0.1$, $Q=1.0$, varying $a=0.1$, $0.3$ and $0.4$ while $\Delta=0$, 1/2 and 1. In this picture, one sees that the slopes of the curves decrease with increasing $\Delta$, whereas the corresponding zeros of $T_{i}$ are not modified by the change of $\Delta$. These curves have approximately the same slopes independent of the values of $a$. On the other hand, increasing $a$ shifts the zeros of $T_{i}$ to smaller values. For example, if $a=0.4$ the zero of $T_{i}$ occurs at $P\approx 0.002$, and for $a=0.1$ it appears at $P\approx 0.001$. In particular, these curves appear in distinct groups depending on the $\Delta$ values, at least for the range chosen here.
@@
Figure 5: Inversion temperature as a function of the pressure for fixed $\omega=-1/6$, $a=0.1$, $Q=1.0$, varying $c$: $c=0.01$ (red line), $c=0.10$ (blue line), $c=1.00$ (black line). Solid lines are obtained with $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
Figure 6: Inversion temperature as a function of the pressure for fixed $\omega=-2/3$, $a=0.1$, $Q=1.0$, varying $c$: $c=0.01$ (red line), $c=0.10$ (blue line), $c=0.15$ (black line). Solid lines are obtained with $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
Figure 7: Inversion temperature as a function of the pressure for fixed $\omega=-1$ (left panel) and $\omega=-4/3$ (right panel), both with $a=0.1$, $Q=1.0$, varying $c$: $c=0.01$ (red line), $c=0.50$ (blue line), $c=1.00$ (black line). Solid lines are obtained with $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$.
@@
Variations of the Kiselev coupling $c$ are presented in Figs. 5, 6, and 7 for the curves $T_{i}\times P$ with fixed $\omega=-1/6$ (Fig. 5), $\omega=-2/3$ (Fig. 6) and $\omega=-1$ (Fig. 7 left panel), $\omega=-4/3$ (Fig. 7 right panel), all of which with $a=0.1$, and $Q=1.0$, for $\Delta=0$, 1/2 and 1. In Fig. 5, the values of $c$ are 0.01, 0.10, and 1.00, while in Fig. 6, the values of $c$ are 0.01, 0.10 and 0.15, and in Fig. 7, the values of $c$ are 0.01, 0.50, and 1.00. In these three pictures, one can see that the increase in the values of $\Delta$ decreases the slope of the inversion temperature curves. The values $\Delta=0$, 1/2 and 1 are apart enough to clearly see the distinction between these groups of curves. In Figs. 5 and 6, there are minimum values for $T_{i}$ that correspond to $P=0$. This is a consequence of the values $\omega=-1/6$ (Fig. 5), and $\omega=-2/3$ (Fig. 6). The increasing values of $c$ and $\Delta$ both imply a decrease in $T_{i}$ for $P=0$. In Fig. 7, the lowest value of $c=0.01$ still implies a finite $T_{i}$ for $P=0$ in both panels, but for $c=0.50$ and $c=1.00$ the minimum value for $T_{i}$ is zero corresponding to non-zero pressures. In this case, the increasing value of $c$ increases the pressure at zero $T_{i}$. In the three pictures, 5, 6, and 7, the curves for fixed $\Delta$ have almost the same slopes and are easily distinguishable from other values of $\Delta$, at least when the range is small. For larger ranges, curves with different $c$ but the same $\Delta$ tend to collapse.

### v2

# Joule-Thomson expansion for quantum corrected AdS-Reissner-Nordström black holes in Kiselev spacetime with Barrow fractal entropy

@@
Joule-Thomson expansion for quantum corrected AdS-Reissner-Nordström black holes in Kiselev spacetime with Barrow fractal entropy
Everton M. C. Abreu1,2 evertonabreu@ufrrj.br    Henrique Boschi-Filho3 boschi@if.ufrj.br    Rafael A. Costa-Silva3 rafaelcosta@pos.if.ufrj.br 1Physics Department, Universidade Federal Rural do Rio de Janeiro, RJ, Brazil.
2Applied Physics Graduate Program, Physics Institute, Universidade Federal do Rio de Janeiro, RJ, Brazil.
3 Instituto de Física, Universidade Federal do Rio de Janeiro, RJ, Brazil.
(9 de março de 2026)
Resumo
Recently, Barrow proposed an extension of the Bekenstein-Hawking black hole entropy to include the effects of fractal geometry through a parameter $\Delta$. Since then, several interesting issues related to this extended entropy have been explored in the literature. In this work, we investigate the effects of the fractal parameter $\Delta$ on the inversion temperature connected to the Joule-Thomson expansion that can be obtained from the thermodynamics of AdS-Reissner-Nordström black holes in Kiselev spacetime. The description of such physical systems involves numerical solutions concerning the fractal parameter $\Delta$. The results are shown by temperature-pressure curves for multiple values of the set of parameters that define the black hole thermodynamics. In conclusion of our analysis, we also show isenthalpic curves corresponding to fixed-mass black hole processes.
@@ I Introduction
The main objective of this paper is to investigate the JT effect modified by the Barrow entropy on quantum corrections and cosmological fluids described by Kiselev spacetime, such as phantom dark energy and quintessence, in an AdS-Reissner-Nordström (AdS-RN) BH. An investigation of JTE in a quantum corrected AdS-Reissner-Nordström black hole in Kiselev spacetime with standard BeH entropy was presented in [59], so the present work can be considered as an extension to study the effects of the fractal Barrow entropy on JTE isenthalpics and inversion temperature in these BHs.
@@ II Joule-Thomson expansion: a brief review
$\mu_{\rm JT}\,=\,\Big(\frac{\partial T}{\partial P}\Big)_{H}\,$, (2)
which is called the JT coefficient. The sign of $\mu_{\rm JT}$ rules the cooling or heating. As pressure decreases, the variation of pressure is negative, but the variation of temperature may be negative or positive. If the temperature variation is negative (positive), then the JT coefficient is positive (negative). Hence, the gas cools (warms).
@@
$\mu_{\rm JT}\,=\,\Big(\frac{\partial T}{\partial P}\Big)_{H}\,=\,\frac{1}{C_{P}}\,\Bigg[T\Big(\frac{\partial V}{\partial T}\Big)_{P}\,-\,V\Bigg]\,.$ (12)
From Eq. (2) we see that the JT coefficient measures the temperature variation of a gas or liquid when it expands without carrying out work keeping the enthalpy constant. At the inversion temperature, there is no cooling or warming and so there is no temperature variation, and the JT coefficient is zero. At $T_{\rm i}^{\rm JT}$ the intermolecular forces are in equilibrium. We will have cooling only at $T$ lower than $T_{\rm i}^{\rm JT}$.
The inversion temperature in the Joule-Thomson expansion is characterized by a zero $\mu_{\rm JT}$, so
$T_{\rm i}^{\rm JT}\,=\,V\,\Big(\frac{\partial T}{\partial V}\Big)_{P}\,\,.$ (13)
@@ III.1 Kiselev black hole
where the metric $d\Omega^{2}\,=\,d\theta^{2}\,+\,\sin^{2}\theta d\phi^{2}$ corresponds to a two-sphere and the horizon function $f(r)$ for a BH with mass $M$, electric charge $Q$, quantum correction $a$, and Kiselev parameters $c$ and $\omega$ is given by
@@ III.2 First law, Barrow entropy, and Inversion Temperature
where the enthalpy $H$ is identified with the BH mass $M=H$, while $\Phi$, $\cal C$, and $\cal A$, respectively, are the potentials of the electric charge, the Kiselev $c$ and the quantum correction parameter $a$. For the inversion temperature, $T_{\rm i}^{\rm JT}$, it can be shown, as we mentioned in the last section, that $\mu_{\rm JT}=0$.
@@
where $\Delta=0$ recover the Bekenstein-Hawking entropy [3, 4] and $\Delta=1$ means that we have the most intricate geometry, the fractal one, i.e., $0\leq\Delta\leq 1$. Hence, we have a fractal event horizon embedded in the Kiselev spacetime. Now we have a quantum correction manifested in the entropy expression. In this Barrow model, it also deforms the structure of the event horizon as a function of the $\Delta$-parameter. In addition, it also deforms the way information is stored in the horizon, creating a quantum microstructure or a fractal one.
@@
Hence, we can write the BH temperature as
$T=\frac{1}{2(2+\Delta)(\sqrt{\pi}r_{\!\!{+}})^{\Delta}}\bigg(8Pr_{\!\!{+}}-\frac{Q^{2}}{\pi r_{\!\!{+}}^{3}}+\frac{1}{\pi\sqrt{r_{\!\!{+}}^{2}-a^{2}}}+\frac{3\omega c}{\pi r_{\!\!{+}}^{2+3\omega}}\bigg)\,$, (21)
which reduces to the Hawking temperature of the AdS-Reissner-Nordström BH in Kiselev spacetime with quantum corrections in the limit $\Delta\to 0$. In addition, the Joule-Thomson inversion temperature is
$T_{\rm i}^{\rm JT}=V\bigg(\frac{\partial T}{\partial V}\bigg)_{P}\,\,$, (22)
@@
$T_{\rm i}^{\rm JT}=\frac{4\pi r_{\!\!{+}}^{3}}{3}\bigg(\frac{\partial T}{\partial r_{\!\!{+}}}\frac{\partial r_{\!\!{+}}}{\partial V}\bigg)_{P}.$ (24)
@@
$T_{\rm i}^{\rm JT}=\frac{1}{6(2+\Delta)(\sqrt{\pi}r_{\!\!{+}})^{\Delta}}\bigg[8P_{i}^{\rm JT}(1-\Delta)r_{\!\!{+}}+\frac{Q^{2}(3+\Delta)}{\pi r_{\!\!{+}}^{3}}-\frac{3\omega c(2+\Delta+3\omega)}{\pi r_{\!\!{+}}^{2+3\omega}}+\frac{\Delta a^{2}-(1+\Delta)r_{\!\!{+}}^{2}}{\pi(r_{\!\!{+}}^{2}-a^{2})^{3/2}}\bigg]\,.$ (25)
Using $T=T_{\rm i}^{\rm JT}$, $P=P_{i}^{\rm JT}$ in Eq. (21), subtracting this equation from Eq. (25), we obtain a relation between $r_{\!\!{+}}$ and $P_{\rm i}^{\rm JT}$, which is given by
$\frac{r_{\!\!{+}}^{3}+r_{\!\!{+}}(3+\Delta)(r_{\!\!{+}}^{2}-a^{2})}{(r_{\!\!{+}}^{2}-a^{2})^{3/2}}+8\,P_{\rm i}^{\rm JT}\,\pi\,r_{\!\!{+}}^{2}(2+\Delta)-\frac{Q^{2}(6+\Delta)}{r_{\!\!{+}}^{2}}+3\,c\,r_{\!\!{+}}^{-1-3\omega}\,\omega\,(5+3\omega+\Delta)=0\,.$ (26)
We will solve this equation numerically to find $r_{\!\!{+}}(P)$. Then, we can substitute this result into Eq. (25) to construct a $T\times P$ plot where we can see the inversion temperature curve. This will be analyzed in detail in the next section.
@@ IV Numerical results
From now on, we will discuss what happens when we compute the variation of $\Delta$ in the face of the other parameters. First, we will consider the case of fixed parameters $a=0.1,\,c=0.1,\,\omega=-1.0,\,Q=1.0$ and we will vary $\Delta$ with a difference $1/4$ in the interval $\Delta\in[0,1]$. Fig. 1 shows the results of this analysis. In the upper panels, we plot the inversion temperature, Eq. (25), against the pressure $P$ in small and large ranges. In these panels, one sees that increasing $\Delta$ reduces the inversion temperature $T_{\rm i}^{\rm JT}$, Eq. (25), for fixed $P$, or equivalently, increases $P$ for fixed $T_{\rm i}^{\rm JT}$. This comes with a change of slope in the curves, without altering the zeros of these functions. Actually, the five curves corresponding to the values of $\Delta$ in the interval $[0,1]$, have the same zero, which is very close to $P_{0}\approx 0.001$. The results of the left panel of Fig. 1 about the decrease of the slope of the inversion temperature curves with increasing values of $\Delta$ are compatible with those found in Fig. 16 of [76] with $F(R)$ gravity, with Figs. 3 and 4 of [77] within Rastall gravity, and with Fig. 4 of [78] within spacetime foam effects.
@@
Then, in the lower panels of Fig. 1, we show the behavior of these functions in the ranges $7\leq P\leq 12$ and $10\leq P\leq 15$, to emphasize the crossing of the inversion temperature curves corresponding to different $\Delta$ values. This means that increasing the fractal dimension $\Delta$ implies decreasing the inversion temperature $T_{\rm i}^{\rm JT}$ for pressures lower than a certain critical value $P_{c}$, and the opposite behavior above this value. In particular, in the lower left panel of Fig. 1, it is seen that the curves for $\Delta=0$ and $\Delta=1$ cross at $P_{c}\approx 10.2$, which can be determined numerically from Eq. (25) using the values $a=0.1,\,c=0.1,\,\omega=-1.0,\,Q=1.0$, for $\Delta=0$ and $\Delta=1$.
@@
Figura 1: Inversion temperature as a function of the pressure for fixed $\omega=-1.0$, $c=0.1$, $a=0.1$, $Q=1.0$, and the corrections to entropy: $\Delta=0$ (blue line), $\Delta=1/4$ (red line), $\Delta=1/2$ (black line), $\Delta=3/4$ (cyan line) and $\Delta=1$ (green line). The upper panels show small and large behaviors of $T_{\rm i}^{\rm JT}\times P$, while the lower panels depict different ranges of these functions, with the purpose of showing the intercept points of these curves for different values of $\Delta$.
@@
Figura 2: Inversion temperature as a function of the pressure for fixed $\omega=-1.0$, $c=0.1$, $a=0.1$, small range, and the values for charge: $Q=1.0$ (green line), $Q=2.0$ (red line), $Q=3.0$ (black line) and $Q=6.0$ (blue line). Solid lines are obtained from $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
In Fig. 2, we show the results for $T_{\rm i}^{\rm JT}\times P$ on both small and large scales, using $\Delta=0$, $\Delta=1/2$, and $\Delta=1$, keeping $a=0.1,\,c=0.1,\,\omega=-1.0$ fixed and varying the charge $Q$ from 1 to 6. Note that increasing $\Delta$ diminishes the slope of the curves $T_{\rm i}^{\rm JT}\times P$, as already found in Fig. 1. In this picture, we also see that the effect on $T_{\rm i}^{\rm JT}$ of increasing the charge $Q$ is similar to the decrease in the fractal parameter of Barrow $\Delta$, so that from the thermodynamical point of view, the charge $Q$ and $\Delta$ have opposite (inverse) behaviors. Note also that the zeros of these curves are different for each charge value $Q$, corresponding to different initial pressures, and we see that these zeros are not modified by changing $\Delta$. For the small range (left panel), one clearly sees a distinction from the curves with $\Delta=0$, $1/2$, and $1$, while in the large range (right panel) these curves get mixed, since for $\Delta=1$, the slope is almost constant, whilst for $\Delta=0$, and $1/2$ they vary significantly, with decreasing slope with increasing pressure. The results of the right panel of Fig. 2 can be compared with those found in Figs. 3 and 4 of [77] within Rastall gravity.
@@
Figura 3: Inversion temperature as a function of the pressure for fixed $Q=2.5$, $c=0.1$, $a=0.1$, varying $\omega$: $\omega=-1.6$ (green line), $\omega=-2/3$ (red line), $\omega=-1$ (black line) and $\omega=-4/3$ (blue line). Solid lines are obtained from $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
In Fig. 3, we plot $T_{\rm i}^{\rm JT}\times P$ for fixed $Q=2.5$, $c=0.1$, $a=0.1$, varying the Kiselev parameter $\omega=-1.6$, 2/3-2/3, 1-1, and 4/3-4/3, for $\Delta=0$, 1/2 and 1. For fixed $\Delta$, the slope of the curves for these values of $\omega$ is very similar, although the zeros of the curves are clearly distinguishable. The slopes of these curves are very sensitive to the variation of $\Delta$, while the zeros are not. Again, increasing $\Delta$ decreases the slope of the curves $T_{\rm i}^{\rm JT}\times P$. In both the small range (left panel) and the large range (right panel), one easily sees a distinction between the curves with different values of $\Delta$, especially in the large range, where the curves with different values of $\omega$ seem to collapse mainly depending on $\Delta$.
@@
Figura 4: Inversion temperature as a function of the pressure for fixed $\omega=-1.0$, $c=0.1$, $Q=1.0$, and the values for the quantum corrections: $a=0.1$ (blue line), $a=0.3$ (red line), $a=0.4$ (black line). Solid lines are obtained with $\Delta=0$, dotted with $\Delta=1/2$, and dashed lines with $\Delta=1$.
@@
To study the effect of the variation of $\Delta$ and the quantum correction parameter $a$, we present in Fig. 4 the graph $T_{\rm i}^{\rm JT}\times P$ for fixed $\omega=-1.0$, $c=0.1$, $Q=1.0$, varying $a=0.1$, $0.3$ and $0.4$ while $\Delta=0$, 1/2 and 1. In this picture, one sees that the slopes of the curves decrease with increasing $\Delta$, whereas the corresponding zeros of $T_{\rm i}^{\rm JT}$ are not modified by the change of $\Delta$. These curves have approximately the same slopes independent of the values of $a$. On the other hand, increasing $a$ shifts the zeros of $T_{\rm i}^{\rm JT}$ to smaller values. For example, if $a=0.4$ the zero of $T_{\rm i}^{\rm JT}$ occurs at $P\approx 0.002$, and for $a=0.1$ it appears at $P\approx 0.001$. In particular, these curves appear in distinct groups depending on the $\Delta$ values, at least for the range chosen here.
@@
Figura 5: Inversion temperature as a function of the pressure for fixed $\omega=-1/6$, $a=0.1$, $Q=1.0$, varying $c$: $c=0.01$ (red line), $c=0.10$ (blue line), $c=1.00$ (black line). Solid lines are obtained with $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
Figura 6: Inversion temperature as a function of the pressure for fixed $\omega=-2/3$, $a=0.1$, $Q=1.0$, varying $c$: $c=0.01$ (red line), $c=0.10$ (blue line), $c=0.15$ (black line). Solid lines are obtained with $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$. Left panel: small range. Right panel: large range.
@@
Figura 7: Inversion temperature as a function of the pressure for fixed $\omega=-1$ (left panel) and $\omega=-4/3$ (right panel), both with $a=0.1$, $Q=1.0$, varying $c$: $c=0.01$ (red line), $c=0.50$ (blue line), $c=1.00$ (black line). Solid lines are obtained with $\Delta=0$, dotted lines from $\Delta=1/2$ and dashed lines from $\Delta=1$.
@@
Variations of the Kiselev coupling $c$ are presented in Figs. 5, 6, and 7 for the curves $T_{\rm i}^{\rm JT}\times P$ with fixed $\omega=-1/6$ (Fig. 5), $\omega=-2/3$ (Fig. 6) and $\omega=-1$ (Fig. 7 left panel), $\omega=-4/3$ (Fig. 7 right panel), all of which with $a=0.1$, and $Q=1.0$, for $\Delta=0$, 1/2 and 1. In Fig. 5, the values of $c$ are 0.01, 0.10, and 1.00, while in Fig. 6, the values of $c$ are 0.01, 0.10 and 0.15, and in Fig. 7, the values of $c$ are 0.01, 0.50, and 1.00. In these three pictures, one can see that the increase in the values of $\Delta$ decreases the slope of the inversion temperature curves. The values $\Delta=0$, 1/2 and 1 are apart enough to clearly see the distinction between these groups of curves. In Figs. 5 and 6, there are minimum values for $T_{\rm i}^{\rm JT}$ that correspond to $P=0$. This is a consequence of the values $\omega=-1/6$ (Fig. 5), and $\omega=-2/3$ (Fig. 6). The increasing values of $c$ and $\Delta$ both imply a decrease in $T_{\rm i}^{\rm JT}$ for $P=0$. In Fig. 7, the lowest value of $c=0.01$ still implies a finite $T_{\rm i}^{\rm JT}$ for $P=0$ in both panels, but for $c=0.50$ and $c=1.00$ the minimum value for $T_{\rm i}^{\rm JT}$ is zero corresponding to non-zero pressures. In this case, the increasing value of $c$ increases the pressure at zero $T_{\rm i}^{\rm JT}$. In the three pictures, 5, 6, and 7, the curves for fixed $\Delta$ have almost the same slopes and are easily distinguishable from other values of $\Delta$, at least when the range is small. For larger ranges, curves with different $c$ but the same $\Delta$ tend to collapse.